40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 8 04:13:12 2012, kiwamu, Update, LSC, update on the locking activity 
    Reply  Fri Mar 9 13:34:13 2012, keiko, Update, LSC, update on the locking activity 
       Reply  Sat Mar 10 02:00:03 2012, keiko, Update, LSC, update on the locking activity 
Message ID: 6393     Entry time: Fri Mar 9 13:34:13 2012     In reply to: 6386     Reply to this: 6398
Author: keiko 
Type: Update 
Category: LSC 
Subject: update on the locking activity 

We tried to measure the sensing matrix for MICH and PRCL last night. They look too much mixed as we expect... the matrix may be posted later. We suspect the IX and IY of the MICH excitation is not balanced very well, although Kiwamu adjusted that about two weeks ago, and it is mixing the dof. We'll try to balance it again, ans see the matrix. 

Keiko, Kiwamu

 

Quote:

[Keiko / Kiwamu]

 Some updates on the locking activity:

  • Started summarizing the data of the Michelson lock in a wiki page:
  • Gradually moving on to the PRMI lock
    • The lock stays for reasonably a long time (~20 min or more)
    • POP22/110 demod signals seemed just ADC noise.
    • A first noise budget is in process
      • The glitches make the noise level worse above 40 Hz or so in both the MICH and PRCL budgets.
    • Sensing matrix will be measured tomorrow
    • The data will be also summarized in a wiki page

 

ELOG V3.1.3-